EQuestion Help X Problem 7-22 (similar to) is$700, and ithas a par vaue of $1,000. The bond has an annual iterest rate of 6% thatspad semarnualy what is the yed to Woldtomatu ty) Assume he market pric e of a 11-year bond for Margaret Inc m matunty of the bond? The yield to meturity of the bond is %. Round to two decimal places.)
Answer #1

Using financial calculator

Input: FV= 1000

PV = -700

N= 11*2 = 22

PMT= 6%*1000/2 = 30

Find I/Y which is the semiannual YTM as 5.35%

Hence YTM = 5.35%*2 = 10.7%

Excel formula rate can also be used to derive the semiannual YTM as


Sign in Book1 - Excel 《Product Activation Failed) AutoSave 수 Share File Insert Page Layout Formulas Data Review View Help Tell me what you want to do Home Wrap Text Percentage 田FF Calibri Conditional Fornat as Cell Insert Delete Fornat FormattingTable Styles- 、. Sort & Find & Editing : Filter Select Paste u . 녀 . 2.A 호 들經垣@werge & Center-5 . % , Styles Cells Number Alignment Clipboard B5 =RATE(22,30,-700,1000) 5.35% 10 12 13 14 16 17 Sheet1 Shet2 + 100 Ready cENG 1:15 AM